An ambulance siren has frequency 250 Hz. The ambulance is headed towards an
accident site with a speed of 90 km/h. Two police officers on separate motor
cycles head for the same accident site: one follows the ambulance with a speed of
80 km/h and the other approaches the accident site from the other direction with a
speed of 80 km/h. What frequency does ambulance siren has for each of the police
officers? Take the speed of sound equal to 340 m/s.
1
Expert's answer
2018-02-14T08:39:07-0500
We use the equation for the Doppler Effect:
f=(1+∆v/c) f_0
1) Police officer follows the ambulance
f=(1+∆v/c) f_0=(1+(22.2- 25)/340)250 Hz = 248 Hz
2) Police officer moving towards the ambulance
f'=(1+∆v'/c) f_0=(1+(25 - 22.2)/340)250 Hz = 252 Hz